# Comprehensive Study of $k$-essence Model: Dynamical System Analysis and Observational Constraints from Latest Type Ia Supernova and BAO Observations

### v2

@@ III Data Analysis
    $\chi^{2}=\sum_{i=1}^{31}\left(\frac{H_{i\ \rm obs}-H_{i\ \rm Model}}{\sigma_{i}}\right)^{2},$ (34)
@@
    $\mu(\tilde{z})=5\,\log_{10}\left[\frac{d_{L}(\tilde{z})}{Mpc}\right]+25,$ (35)
@@
    $d_{L}(\tilde{z})=c(1+\tilde{z})\int_{0}^{\tilde{z}}\frac{dz^{\prime}}{H(z^{\prime})},$ (36)
@@
    $\chi^{2}_{\text{SNe Ia}}=\Delta\vec{D}^{T}C^{-1}_{\rm stat+syst}\Delta\vec{D},$ (37)
@@
    $\Delta\vec{D}=\mu(\tilde{z}_{i})-\mu_{th}(\tilde{z}_{i},H_{0},\Omega_{m_{0}},y_{0}).$ (38)
@@
    $\chi^{2}_{\rm BAO}=\sum_{i=1}^{N}\left(\frac{X_{i}}{\sigma_{i}}\right)^{2}$ (39)
@@
    $X_{i}=\text{Data}_{i}-\text{Model}_{i}\ ,$ (40)
@@
    $D_{V}(\tilde{z})\equiv\left({\tilde{z}\ D_{M}^{2}(\tilde{z})D_{H}}\right)^{1/3}\ .$ (41)
@@
    $c_{s}(\tilde{z})=\frac{c}{\sqrt{3(1+\frac{3\rho_{B}}{4\rho_{\gamma}})}}\,.$ (43)
@@
$P(\theta|D,M)=\frac{P(D|\theta,M)P(\theta|M)}{P(D|M)},$ (44)
@@
$P(D|\theta,M)\equiv\exp\left(\frac{-\chi^{2}(\theta)}{2}\right).$ (45)
@@
Marginalized 1D and 2D posterior distribution of the dynamical variables and model parameters for the $\Lambda$CDM model, Model I and Model II for various data sets are shown in Fig. [7]. The mean and standard deviation of the distribution are considered as the best-fit value and error corresponding to each model parameters, which are summarized in Tabs. [3, 4, 5] respectively. In the table, we provided the dynamical variables and model parameters corresponding to Model I and Model II and specified the corresponding prior ranges. These prior ranges were selected based on the analysis conducted in the previous dynamical stability section. Note that the parameter ranges specified in Eqs. (18) and (33) are extended in Tabs. [4, 5].
For Model I, which refers to the $k$-essence field with inverse square potential, there are mainly two primary variables, $(x,y)$, whose initial conditions $(x_{0},y_{0})$ must be varied. However, in the table, we have listed $x_{0}$ and $\Omega_{m0}$ instead of $y_{0}$. The reason for this is clarified in Eq. (21). On the other hand, for Model II, we have followed the same approach, constraining all the necessary parameters except $\mathcal{M}_{0}$. The parameter $\mathcal{M}_{0}$ is fixed to a certain value $\mathcal{M}_{0}=8.0$, as this parameter must be a positive number. We emphasize that although this parameter is fixed, the field variables $(x_{0},y_{0})$ are still varying, as can be seen from the Hubble equation Eq. (27), which reads as:
$y_{0}=\frac{\log\left(-\frac{12(\Omega_{m0}-1)}{\mathcal{M}_{0}\left(3x^{4}-2x^{2}\right)}\right)}{\beta}.$ (48)
@@
Additionally, we evaluated the information criteria, specifically AIC and BIC, for the models. For this comparison, $\Lambda$CDM was considered the base model. The model having lower value of AIC or BIC is preferred over the other. From Tab. [3], [4] and [5], it is evident from the AIC criterion that Model 1 and Model 2 are preferred over the $\Lambda$CDM model for most data combinations. However, according to the BIC, the $\Lambda$CDM model is preferred over Model I and Model II for all data combinations, as BIC more effectively penalizes models with extra number of parameters.
Before we end this section we want to opine on the pure kinetic $k$-essence models, which are thought to be simpler to handle as there are no dynamic potential terms and where the Lagrangian density of the $k$-essence field is purely a function of $X$. In these models $V$ is a constant parameter. Such a case is discussed in ref. [36], where the author assumes that the Lagrangian of the $k$-essence field has an extrema at some point $X=X_{0}$. It is seen that if one Taylor expands the Lagrangian at $X=X_{0}$ then one obtains a theory where the energy density of the $k$-essence field near $X_{0}$ can be represented as a sum of two separate terms. One of these terms can be interpreted to be coming from a dark energy like component and the other term scales with the scale-factor as $a^{-3}$, specifying a dark matter like component. Pure kinetic $k$-essence fields can in a certain way unify dark matter and dark energy sectors. The resultant theory has some parameters whose values require to be ascertained by observational data. We see that with our present sets of data it is practically impossible to give any bound on the parameters appearing in the theory. Although a $k$-essence sector with a non trivial potential can be tackled by our method and we can properly give bounds on various parameters and initial conditions appearing in the theory, it is pratically impossible to do such an analysis with a less complicated model. One of the reasons why this happens is perhaps related to the high level of generality of the problem discussed in ref. [36]. This theory is very general in nature and does not specify any particular form of the Lagrangian of the purely kinetic $k$-essence. Due to this general feature the energy density of the field near the extrema can be expressed in terms of six constants ($V_{0}$, $X_{0}$, $F_{0}$, $F_{2}$, $a_{1}$ and $\epsilon_{1}$). It becomes difficult to put reasonable bounds on all these six parameters using the present set of observational data. As we do not get any positive result for this case we do not present any more detail about this theory here. Readers who want to get a feeling of the difficulty one faces, if one wants to compare it with data, can consult Appendix A where the problem is briefly discussed.
IV Conclusion
@@
Many studies have been conducted with the inverse square potential, but few have analyzed the exponential type potential with this particular form of the Lagrangian. One key challenge in the dynamics of the exponential potential is the need for more than two dynamical variables to close the autonomous system of equations, resulting in a 3D phase space. Due to the different structure of dynamical variables compared to the inverse square potential, the exponential model does not yield critical points that signify the current accelerating or phantom phase. The critical points only describe either far past or future epochs. Thus, model parameters such as $\beta$ and $\mathcal{M}_{0}$ could not be constrained with the arguments presented for Model I. Therefore, we relied on numerical evolution and skipped the critical point analysis, varying the initial conditions and model parameters within a specific range to see whether the model produces different phases of the universe. We found that for certain ranges, the model produces a matter-dominated phase, and as the system enters the current epoch, the system’s equation of state becomes lower than the background fluid’s EoS, exhibiting tracking behavior. Thus, the exponential model can produce a stable accelerating (phantom) phase followed by an attractive matter phase for certain ranges, alleviating the fine-tuning issue.
@@
We further assessed the viability of the models using statistical criteria, specifically the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC), and compared the results with those of the $\Lambda$CDM model. For most of the combined datasets, both models outperformed $\Lambda$CDM according to AIC criterion. However, $\Lambda$CDM remained more favorable for all the data combinations when we considered BIC criterion. This is because BIC criterion penalizes models with additional parameters. This suggests that while the current models provide good fit to the all the data combinations, the $\Lambda$CDM model is preffered as compared to k-essence models.
Acknowledgements.
@@
Appendix A A note on pure kinetic $k$-essence

### v3

@@ 3 Data Analysis
    $\chi^{2}=\sum_{i=1}^{31}\left(\frac{H_{i\ \rm obs}-H_{i\ \rm Model}}{\sigma_{i}}\right)^{2},$ (3.1)
@@
    $\mu(\tilde{z})=5\,\log_{10}\left[\frac{d_{L}(\tilde{z})}{Mpc}\right]+25,$ (3.2)
@@
    $d_{L}(\tilde{z})=c(1+\tilde{z})\int_{0}^{\tilde{z}}\frac{dz^{\prime}}{H(z^{\prime})},$ (3.3)
@@
    $\chi^{2}_{\text{SNe Ia}}=\Delta\vec{D}^{T}C^{-1}_{\rm stat+syst}\Delta\vec{D},$ (3.4)
@@
    $\Delta\vec{D}=\mu(\tilde{z}_{i})-\mu_{th}(\tilde{z}_{i},H_{0},\Omega_{m_{0}},y_{0}).$ (3.5)
@@
    $\chi^{2}_{\rm BAO}=\sum_{i=1}^{N}\left(\frac{X_{i}}{\sigma_{i}}\right)^{2}$ (3.6)
@@
    $X_{i}=\text{Data}_{i}-\text{Model}_{i}\ ,$ (3.7)
@@
    $D_{V}(\tilde{z})\equiv\left({\tilde{z}\ D_{M}^{2}(\tilde{z})D_{H}}\right)^{1/3}\ .$ (3.8)
@@
    $c_{s}(\tilde{z})=\frac{c}{\sqrt{3(1+\frac{3\rho_{B}}{4\rho_{\gamma}})}}\,.$ (3.10)
@@
$P(\theta|D,M)=\frac{P(D|\theta,M)P(\theta|M)}{P(D|M)},$ (3.11)
@@
$P(D|\theta,M)\equiv\exp\left(\frac{-\chi^{2}(\theta)}{2}\right).$ (3.12)
@@
Marginalized 1D and 2D posterior distribution of the dynamical variables and model parameters for the $\Lambda$CDM model, Model I and Model II for various data sets are shown in Fig. [7]. The mean and standard deviation of the distribution are considered as the best-fit value and error corresponding to each model parameters, which are summarized in Tabs. [4, 5, 6] respectively. In the table, we provided the dynamical variables and model parameters corresponding to Model I and Model II and specified the corresponding prior ranges. These prior ranges were selected based on the analysis conducted in the previous dynamical stability section. Note that the parameter ranges specified in Eqs. (2.18) and (2.33) are extended in Tabs. [5, 6].
For Model I, which refers to the $k$-essence field with inverse square potential, there are mainly two primary variables, $(x,y)$, whose initial conditions $(x_{0},y_{0})$ must be varied. However, in the table, we have listed $x_{0}$ and $\Omega_{m0}$ instead of $y_{0}$. The reason for this is clarified in Eq. (2.21). On the other hand, for Model II, we have followed the same approach, constraining all the necessary parameters except $\mathcal{M}_{0}$. The parameter $\mathcal{M}_{0}$ is fixed to a certain value $\mathcal{M}_{0}=8.0$, as this parameter must be a positive number. We emphasize that although this parameter is fixed, the field variables $(x_{0},y_{0})$ are still varying, as can be seen from the Hubble equation Eq. (2.27), which reads as:
$y_{0}=\frac{\log\left(-\frac{12(\Omega_{m0}-1)}{\mathcal{M}_{0}\left(3x^{4}-2x^{2}\right)}\right)}{\beta}.$ (3.15)
@@
Additionally, we evaluated the information criteria, specifically AIC and BIC, for the models. For this comparison, $\Lambda$CDM was considered the base model. The model having lower value of AIC or BIC is preferred over the other. From Tab. [4], [5] and [6], it is evident from the AIC criterion that Model I and Model II are preferred over the $\Lambda$CDM model for most data combinations. However, according to the BIC, the $\Lambda$CDM model is preferred over Model I and Model II for all data combinations, as BIC more effectively penalizes models with extra number of parameters.
Before we end this section we want to opine on the pure kinetic $k$-essence models, which are thought to be simpler to handle as there are no dynamic potential terms and where the Lagrangian density of the $k$-essence field is purely a function of $X$. In these models $V$ is a constant parameter. Such a case is discussed in ref. [36], where the author assumes that the Lagrangian of the $k$-essence field has an extrema at some point $X=X_{0}$. It is seen that if one Taylor expands the Lagrangian at $X=X_{0}$ then one obtains a theory where the energy density of the $k$-essence field near $X_{0}$ can be represented as a sum of two separate terms. One of these terms can be interpreted to be coming from a dark energy like component and the other term scales with the scale-factor as $a^{-3}$, specifying a dark matter like component. Pure kinetic $k$-essence fields can in a certain way unify dark matter and dark energy sectors. The resultant theory has some parameters whose values require to be ascertained by observational data. We see that with our present sets of data it is practically impossible to give any bound on the parameters appearing in the theory. Although a $k$-essence sector with a non trivial potential can be tackled by our method and we can properly give bounds on various parameters and initial conditions appearing in the theory, it is pratically impossible to do such an analysis with a less complicated model. One of the reasons why this happens is perhaps related to the high level of generality of the problem discussed in ref. [36]. This theory is very general in nature and does not specify any particular form of the Lagrangian of the purely kinetic $k$-essence. Due to this general feature the energy density of the field near the extrema can be expressed in terms of six constants ($V_{0}$, $X_{0}$, $F_{0}$, $F_{2}$, $a_{1}$ and $\epsilon_{1}$). It becomes difficult to put reasonable bounds on all these six parameters using the present set of observational data. As we do not get any positive result for this case we do not present any more detail about this theory here. Readers who want to get a feeling of the difficulty one faces, if one wants to compare it with data, can consult Appendix C where the problem is briefly discussed.
4 Conclusion
@@
Many studies have been conducted with the inverse square potential, but few have analyzed the exponential type potential with this particular form of the Lagrangian. One key challenge in the dynamics of the exponential potential is the need for more than two dynamical variables to close the autonomous system of equations, resulting in a 3D phase space. Due to the different structure of dynamical variables compared to the inverse square potential, the exponential model does not yield critical points that signify the current accelerating or phantom phase. The critical points only describe either far past or future epochs. Thus, model parameters such as $\beta$ and $\mathcal{M}_{0}$ could not be constrained with the arguments presented for Model I. Therefore, we relied on numerical evolution by varying the initial conditions and model parameters within a specific range to see whether the model produces different phases of the universe. We found that for certain ranges, the model produces a matter-dominated phase, and as the system enters the current epoch, the system’s equation of state becomes lower than the background fluid’s EoS, exhibiting tracking behavior. Thus, the exponential model can produce a stable accelerating (phantom) phase followed by the matter phase for certain ranges, alleviating the fine-tuning issue.
@@
We further assessed the viability of the models using statistical criteria, specifically the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC), and compared the results with those of the $\Lambda$CDM model. For most of the combined datasets, both models outperformed $\Lambda$CDM according to AIC criterion. However, $\Lambda$CDM remained more favorable for all the data combinations when we considered BIC criterion. This is because BIC criterion penalizes models with additional parameters. This suggests that while the current models provide good fit to the all the data combinations, the $\Lambda$CDM model is preferred as compared to $k$-essence models.
However, recent studies indicate that we require a cosmological model that enhance the late-time cosmic acceleration to alleviate the Hubble tension. Phantom dark energy models, characterized by an equation of state parameter $w$ evolving just below $-1$, could potentially resolve the Hubble tension [85, 86, 87, 88]. The increase in acceleration within the phantom dark energy models raises the best-fit value of $H_{0}$ as compared to the $\Lambda$CDM model, brings it closer to the values derived from local distance ladder measurements. The $k$-essence models considered in this study showing phantom nature can potentially alleviate the Hubble tension. However, achieving an $H_{0}$ value from the CMB data that aligns with local measurements is insufficient on its own; a superior statistical fit of phantom models relative to the standard $\Lambda$CDM model and dynamical dark energy models is also required. Further analysis, including CMB data, is crucial to determine whether the $k$-essence models produce an $H_{0}$ value that is more consistent with the SH0ES measurement and to determine whether the $k$-essence models produce superior fit to CMB data as compared to the $\Lambda$CDM.
Before we conclude, it is important to note that although the combined data sets marginally favored the phantom nature of the model, phantom phases of accelerated expansion are not physically well motivated due to their ability of producing a diverging phantom energy density in finite cosmological time [89]. Except this there can be other difficulties in phantom cosmology [90]. In our analysis, for both models (corresponding to different potentials), the universe predominantly exhibits non-phantom behavior in the late-time phase, except for the dataset combination CC+BAO+DESBAO+PP+DES. This fact shows that our work is not unduly biased for phantom like behavior of the $k$-essence sector but as $k$-essence theories allow phantom like phases, our result to a certain extent, depends on phantom like behavior of the $k$-essence field. We can conclude by saying that alternative approaches beyond phantom dark energy models, along with additional data points at intermediate redshifts, are essential to effectively address the Hubble tension problem.
Acknowledgments
@@
Appendix A Critical Points at infinity in the case of power law potential
In this appendix, we derive the critical points for the power-law potential by transforming the predefined variables. The ranges of the predefined variables are:
$-\infty<x<+\infty,\quad 0<y<+\infty,\quad 0\leq\Omega_{m}<1\ .$ (A.1)
Using these variables, the Hubble constraint is expressed as:
$y^{2}=1-\Omega_{m}+x^{2}\ .$ (A.2)
Since $\Omega_{m}$ is restricted to a finite range, for very large $x\to\infty$, it follows that $y$ also tends to infinity. To address this asymptotic behavior, where both $x$ and $y$ diverge, we introduce a transformation of variables using hyperbolic trigonometric functions:
$x=u\sinh(v),\quad y=u\cosh(v)\,.$ (A.3)
In these new variables, the Hubble constraint becomes:
$1=u^{2}+\Omega_{m}\ .$ (A.4)
Here, the fractional energy density of the field is given by $\Omega_{\phi}=u^{2}$, which is constrained within the range $0\leq u^{2}\leq 1$. Thus, the variable $u$ is now restricted to a finite range.
The other variable, $v$, is related to $x$ and $y$ as follows:
$v=\tanh^{-1}(x/y)\ .$ (A.5)
The range of $v$ remains:
$-\infty<v<+\infty\ .$ (A.6)
To compactify $v$ within a finite range, allowing for the identification of critical points at infinity, we apply the following Poincaré transformation:
$\bar{v}=\frac{v}{\sqrt{1+v^{2}}}\ .$ (A.7)
This maps $\bar{v}$ into the range:
$-1<\bar{v}<+1\ .$ (A.8)
Using the transformed variables $(u,\bar{v})$, the dynamics of the system are now confined to finite ranges. The corresponding autonomous system of equations becomes:
$u^{\prime}$ $=$ $-\frac{1}{2}u\left(u^{2}-1\right)\left(\cosh\left(\frac{2\bar{v}}{\sqrt{1-\bar{v}^{2}}}\right)-2\right),$ (A.9)
$\bar{v}^{\prime}$ $=$ $-\frac{\left(1-\bar{v}^{2}\right)^{3/2}\coth\left(\frac{\bar{v}}{\sqrt{1-\bar{v}^{2}}}\right)\left(\sqrt{6}\lambda u\sinh\left(\frac{\bar{v}}{\sqrt{1-\bar{v}^{2}}}\right)+\cosh\left(\frac{2\bar{v}}{\sqrt{1-\bar{v}^{2}}}\right)-5\right)}{4\coth^{2}\left(\frac{\bar{v}}{\sqrt{1-\bar{v}^{2}}}\right)-2}.$ (A.10)
Points $u$ $\bar{v}$ $\Omega_{\phi}$ $\Omega_{m}$ $\omega_{\rm eff}$ Stability
$P_{1}$ 0 0.753 0 1 0 Saddle
$P_{2}$ 1 $\frac{\cosh^{-1}\left(0.5\sqrt{3.\lambda^{2}-1.\sqrt{9.\lambda^{4}+48.\lambda^{2}}+12.}\right)}{\sqrt{\cosh^{-1}\left(0.5\sqrt{3.\lambda^{2}-1.\sqrt{9.\lambda^{4}+48.\lambda^{2}}+12.}\right)^{2}+1.}}$ 1 0 Fig. [8a] Fig. [8b]
$P_{3}$ 1 $\frac{\cosh^{-1}\left(0.5\sqrt{3.\lambda^{2}+\sqrt{9.\lambda^{4}+48.\lambda^{2}}+12.}\right)}{\sqrt{\cosh^{-1}\left(0.5\sqrt{3.\lambda^{2}+\sqrt{9.\lambda^{4}+48.\lambda^{2}}+12.}\right)^{2}+1.}}$ 1 0 Fig. [8c] Fig. [8d]
$P_{4}$ 1 0 1 0 1/3 Unstable
$P_{5}$ 0 1 0 1 0 Unstable
$P_{6}$ $\frac{\sqrt{3}}{\lambda}$ $0.55$ $3/\lambda^{2}$ $1-\frac{3.}{\lambda^{2}}$ 0 Stable ($\lambda>\sqrt{3}$)
Table 7: Critical points for the compactified variables $(u,\bar{v})$ in the context of the power-law potential.
Here, the prime denotes differentiation with respect to the logarithmic time variable $dN=d\log(a)$. The autonomous system reveals invariant sub-manifolds at $u=0$ and $v=0$ lines. These invariant manifolds ensure that trajectories originating in the $u>0,\bar{v}>0$ quadrant remain confined within it.
[…]
  • Points $P_{1,5}$: Both points correspond to fluid-dominated energy density with a vanishing effective EoS, indicative of a matter-dominated phase. However, $P_{1}$ exhibits a saddle nature, while $P_{5}$ is unstable.
  • Point $P_{4}$: This point is characterized by a field-dominated energy density and an effective EoS of $1/3$, representing non-thermal radiation. This point is unstable.
  • Points $P_{2,3}$: These points depend on the parameter $\lambda$ and correspond to field-dominated phases. $P_{2}$ leads to an accelerating solution, while $P_{3}$ results in a phantom phase. Figures [8a] and [8c] depict the variation of the effective EoS with $\lambda$, while Figures [8b] and [8d] show the eigenvalue analysis, highlighting their stability.
  • Point $P_{6}$: This point exists only for $\lambda\geq\sqrt{3}$. While it represents a stable matter-dominated phase, it is not physically viable, as the universe cannot remain stabilized in the matter-dominated epoch.
Overall, this analysis reveals a dynamical behavior similar to the original variables, with the additional presence of critical points at infinity. However, these points do not affect the late-time dynamics since they are non-accelerating.
The overall behavior of the system becomes evident from the phase-space plot in Figure [9] for different values of $\lambda$. For specific $\lambda$, the point $P_{6}$ violates the Hubble constraint, rendering it invisible in the phase space. The gray-shaded region represents the allowed domain defined by $10\leq\Omega_{\phi}<$ 1. The phase space is divided into regions with distinct dynamics: The green region corresponds to the accelerating regime ($-1\leq\omega_{\rm eff}\leq-1/3$). The red region indicates the phantom regime ($\omega_{\rm eff}<-1$). The yellow region represents the radiation-dominated phase ($0.310.29<\omega_{\rm eff}<$ 0.31). The phase-space trajectories originating from $P_{4}$ initially converge towards $P_{1}$, corresponding to the matter-dominated phase. Since $P_{1}$ is a saddle point, these trajectories are repelled and eventually stabilize at the accelerating attractor $P_{2}$ (for $\lambda=0.2$). Similarly, trajectories from $P_{5}$ also stabilize at $P_{2}$. However, a few trajectories bypass move directly toward the stable phantom point $P_{3}$.
For $\lambda=-0.2$, the trajectories originating from $P_{4}$ also stabilize at the phantom point $P_{3}$. For $\lambda>1$, the trajectories stabilize at $P_{2}$ or $P_{3}$, but neither of these points lies in the accelerating or phantom regions, making the dynamics irrelevant for this case.
It is seen that only point $P_{4}$ has one of the coordinates $\bar{v}$ to be 0 designating that it corresponds to the point at where both $x$ and $y$ tend to $\infty$. We did not discuss about this critical point in our main analysis given in the text as we were mainly interested in the critical points in the finite regions of phase space. For $\bar{v}=0$ we must have both $x$ and $y$ tends to infinity (but $y$ tends to infinity at a much faster rate than $x$) such that $x/y$ tends to zero. This point does not play a prominent role in the late time universe and consequently only the critical points in the finite region of phase space remain relevant for our purpose  [^1] One must note that the point with $\bar{v}=1$ does not specify any of our old variables $x$ or $y$ are infinity. Although this point designates $v\to\infty$, but $x$ and $y$ are finite for this value of $v$. In summary, the compactified phase space provides a richer dynamical perspective compared to the original variables. Nevertheless, the overall behavior remains consistent, except for the existence of a critical point at infinity in the compactified variables. Importantly, these points do not influence the dynamics at the current epoch, justifying the use of the original variables for the data analysis presented in Sec. 3.
Figure 8: Variation of the effective equation of state (EoS) and eigenvalues for the power-law potential in compactified variables $(u,\bar{v})$: (a) and (b) correspond to point $P_{2}$, while (c) and (d) correspond to point $P_{3}$.
Figure 9: Phase space for the power-law potential using the compactified variables $(u,\bar{v})$, shown for different values of $\lambda$.
Appendix B A compact phase space analysis for the case of exponential potential
In this section, we discuss the compactification of dynamical variables for an exponential type potential. The predefined ranges for these variables are:
$-\infty<x<+\infty,\quad-\infty<y<+\infty,\quad-1<z<1.$ (B.1)
To compactify $y$, we apply the Poincaré transformation, defining a new variable $\bar{y}$ as:
$\bar{y}=\frac{y}{\sqrt{1+y^{2}}},$ (B.2)
with this transformation, $\bar{y}$ is now constrained within:
$-1<\bar{y}<+1.$ (B.3)
In these transformed variables, the Hubble constraint equation becomes:
$1=\Omega_{m}+\frac{\mathcal{M}_{0}\left(\frac{3x^{4}}{4}-\frac{x^{2}}{2}\right)(1-z)^{2}e^{\frac{\beta\bar{y}}{\sqrt{1-\bar{y}^{2}}}}}{3(z+1)^{2}}.$ (B.4)
It is evident that by compactifying $y$ into $\bar{y}$, the variable $x$ is also compactified. The autonomous system of equations now becomes:
$x^{\prime}$ $=$ $\frac{-3x^{3}+\frac{\beta\left(\frac{x^{2}}{2}-\frac{3x^{4}}{4}\right)(1-z)}{z+1}+3x}{3x^{2}-1},$ (B.5)
$\bar{y}^{\prime}$ $=$ $\frac{x(1-z)}{z+1}\left(1-\bar{y}^{2}\right)^{3/2},$ (B.6)
$z^{\prime}$ $=$ $-\frac{3((1-z)(z+1))\left(\frac{1}{3}\mathcal{M}_{0}\left(\frac{x^{4}}{4}-\frac{x^{2}}{2}\right)\left(\frac{1-z}{z+1}\right)^{2}\exp\left(\beta\frac{\bar{y}}{\sqrt{1-\bar{y}^{2}}}\right)+1\right)}{4}\,.$ (B.7)
With this transformation, we encounter a similar issue (as faced previously when dealing with the equations in the main text): the autonomous system diverges as $z\to-1$. To address this, we redefine the time variable as $dN\to(1+z)d\tilde{N}$. The resulting equations can be written down in a straightforward manner although we have not written them here. A potential concern is the behavior of the exponential argument in $z^{\prime}$ as $\bar{y}\to 1$, which might cause $z^{\prime}$ to diverge. However, the structure of $z^{\prime}$ includes the factor $(1-z)(1+z)$, so when $\bar{y}$ approaches 1, $z$ simultaneously approaches either ±$\pm 1$, keeping $z^{\prime}$ finite. Thus, the autonomous system remains differentiable, and $\bar{y}\to\pm 1$ (or equivalently, $y\to\mp\infty$) remains a valid limit. This behavior is illustrated in Fig. [4], for the non-transformed variable evolution.
In this case, the critical points listed in Tab. [8] are identical to those for the non-transformed variables (as presented in the main text), so further discussion on them is unnecessary.
Points $(x,\bar{y},z)$ $\Omega_{\phi}$ $\Omega_{m}$ $\omega_{\rm eff}$
$P_{1,2}$ $(0,\text{Any},\mp 1)$ 0 1 0
$P_{3,4}$ $(\mp 1,\text{Any},1)$ 0 1 0
Table 8: The nature of critical points corresponding to an exponential potential with compacted $\bar{y}$.
It can be seen that for all the critical points in this case ($H\to\infty$ or $H\to 0$), the effective EoS of the system is zero and consequently these critical points are not important for producing accelerated expansion of the universe. Moreover as far as comparison with data is concerned we get important results, which matches with our expectations about the present day universe, showing that a probable future or past asymptotic fixed point with unnatural values of the Hubble parameter, does not affect our predictions. In such a case a critical point analysis does not serve our purpose.
In our case we work out the dynamics using numerical simulations varying initial conditions and model parameters within the ranges:
$-0.1<\bar{y}_{0}<0.4,\quad z_{0}=0,\quad 6<\mathcal{M}_{0}<8,\quad-0.1<\beta<0.4,\quad\Omega_{\rm m0}=0.3\,.$ (B.8)
The initial condition for $x$ can be determined from the constraint relation. The numerical evolution is shown in Fig. [10]. In the past epoch, $x$ takes larger values, but the fractional energy densities for both the field and fluid remain finite and within physically acceptable limits. In the future epoch, $\bar{y}$ approaches $1$, where field energy density dominates, and matter density vanishes, signifying a dark energy-dominated behavior with an effective EoS transitioning between accelerating and phantom regimes. This indicates that a $k$-essence field with an exponential potential can produce stable accelerating or phantom solutions at late times, with behavior similar to that shown in Fig. [4].
Figure 10: Numerical evolution of the system with an exponential potential, using the compactified variable $\bar{y}$.
Appendix C A note on pure kinetic $k$-essence